OpenStudy (anonymous):

Can someone help me with this?? solve for y: (y-8)-(y+9) = 2y

OpenStudy (anonymous):

open the brackets first..

OpenStudy (anonymous):

ok y-8-y+9=2y

OpenStudy (anonymous):

??

OpenStudy (anonymous):

look negative sign outside bracket means you have to change all the signs like y-8-y-9=2y

OpenStudy (anonymous):

oh ok, now do I combine like terms?

OpenStudy (anonymous):

yes now combine like terms...

OpenStudy (anonymous):

y=-17

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

let me solve this...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y-8-y-9=2y y-y -8 -9 = 2y -17 = 2y divide both Sides by "2" -17/2 = 2y/2 -17/2 = y

OpenStudy (anonymous):

@WhyteMonks got it now..??

OpenStudy (anonymous):

oh ok.. thanks

OpenStudy (anonymous):

no problem... :)